3. The following is information from the national income accounts for a hypothetical country:
GDP $6,000
Gross investment 800
Net investment 200
Consumption 4,000
Government purchases of goods and services 1,100
Government budget surplus 30
What is
a. NDP? d. Disposable personal income?
b. Net exports? e. Personal saving?
c. Government taxes minus transfers?
1
Expert's answer
2021-08-09T15:45:20-0400

GDP: Y = Rs. 6000,

Gross Investment I = 800,

Net Investment NI = 200,

Consumption C = 4000,

Govt. purchases of goods & services G = 1100,

Govt. Budget Surplus BS = T - G = 30.

a) Depreciation = I - NI = 800 - 200 = 600.

NDP = Y - depreciation = 6000 - 600 = 5400.

b) Net exports NX = Y - C - I - G = 6000 - 4000 - 800 - 1100 = 100.

c) Govt. taxes minus transfers T = G + BS = 1100 + 30 = 1130.

c) Disposable personal income Yd = Y - T = 6000 - 1130 = 4870.

e) Personal Saving S = Yd - C = 4870 - 4000 = 870.